OpenStudy (anonymous):

Please help me x^2-5x-y-1=2x^2+y-9 find y??

OpenStudy (cwrw238):

you aim to get y only on one side of the '=' first add y to both sides of the equation x^2-5x-y-1=2x^2+y-9 +y +y this gives x^2 - 5x - 1 = 2x^2 +2y - 9

OpenStudy (cwrw238):

now by subtracting 2x^2 and adding 9 to both sides you'll get 2y on its own on the right hand side

OpenStudy (cwrw238):

can u do that?

OpenStudy (anonymous):

yeah i can

OpenStudy (cwrw238):

x^2 - 5x - 1 = 2x^2 +2y - 9 - 2x^2 +9 = -2x^2 +9 gives

OpenStudy (cwrw238):

can you write the next line?

OpenStudy (anonymous):

then wats next?

OpenStudy (cwrw238):

you need to simplify x^2 - 2x^2 and -1 + 9

OpenStudy (cwrw238):

after you have done the above you divide both sides by 2 ( because you have 2y and you need y)

OpenStudy (anonymous):

2y=3x^2-5x+10?

OpenStudy (cwrw238):

no 2y = -x^2 - 5x + 8